OpenStudy (anonymous):

pleaseFind the domain of f(x)=x+sqroot (25-(x^2))

OpenStudy (anonymous):

For the square root to return real values, we need 25-x^2 >= 0 so, 25>=x^2 x^2<=25 So: -5<=x<=5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

note that: 25-x^2>=0 means that: (5-x)(5+X)>=0 So, looking at a number line at (5-x) and (5+x) to see when each of them is positive , negative or zero . Then looking at their product (5-x)(5+x) ; is how we see that: 25-x^2 >= 0 when -5 <= x <= 5 <------------+-----------------------+---------------> -5 5 (5-x) + + + + + + + + + + + + + + + + + + + 0 - - - - - - - - - - - (5+x) - - - - - - 0 + + + + + + + + + + + + + + + + + + + + + + + (5-x)(5+x) - - - - - - - 0 + + + + + + + + + + + + + 0 - - - - - - - - - - -

OpenStudy (anonymous):

oh i see thanks a lot again :)